Colby worked three more hours on Tuesday than he did on Monday. On Wednesday, he worked one hour more than twice the number of hours that he worked on Monday. If the total number of hours is two more than five times the number of hours worked on Monday, how many hours did he work on Monday?
Answer by Lyd999(3)

I'm going to use Monday as X
Monday = X
Tuesday = X + 3
(Problem says three more hours than he worked on Monday)
Wednesday = 2x + 1
(Problem says he worked 1 hour more than two times the number of hours on monday)
This side of the equation would be -
X + (X +3) + (2X + 1)
That's monday plus tuesday plus wednesday.
Then you would set up the other side of the equation.
This would be the total number of hours so it would be equal to monday, tuesday and wednesday combined.
2 + (5x)
Two hours more than five times the amount of Monday (X)

Now we put this together to have an equation
X + (X + 3) + (2x+ 1) = 2 + 5x
Now we need to collect like terms
4x + 4 = 2 + 5x
I just simplified the left side of the equation
Now I will subtract 4x from the left side to get all the variables on one side
4 = 2 + x
Now I subtract 2 to get the numbers both on one side
2 = x
So, Colby worked 2 hours on Monday.

I hoped this helped!